The length of a rectangle is 8cm more than four times the width. If the perimeter of the rectangle is 56cm what are the dimensions?
Answer by TheInstructor(29)

Maybe I should create a lesson on this, nevertheless:

Let's break it down. Okay so your length is 8cm more than width. How can we re-write this into the form of an equation?

Let your x = Width
Simply put: L = 4x + 8

Okay your perimeter is given as 56cm. Now we know that Perimeter = 2Length + 2Width right (2L+2W)?

So: 2(4x+8)+ 2(x) = 56
Therefore: 8x+16 + 2x = 56
10x + 16 = 56
10x = 40

SO YOUR X IS THEN EQUAL TO 4CM. Now remember your length is 4x+8. So sub your x=4, into the length equation to give you a dimension of 24cm.

So as a result;
Length: 24cm
Width: 4cm
If you want to check it sub your dimensions into the 2L+2W formula; which should give you an answer of 56.

Hope that helped! :)